OpenStudy (roseflight):

What is the limit of sin(2x)/(3x) as x approaches 0

OpenStudy (anonymous):

Would you like the step by step?

OpenStudy (roseflight):

Yes please

OpenStudy (turingtest):

\[\frac{\sin(2x)}{3x}=\frac13\cdot\frac{\sin(2x)}x=\frac23\cdot\frac{\sin(2x)}{2x}\]

OpenStudy (roseflight):

Sorry, but I still don't get how that would equal 2/3.

OpenStudy (turingtest):

did you know that\[\lim_{x\to0}\frac{\sin x}x=1\]?

OpenStudy (roseflight):

Haha I didn't remember that for some reason. Thanks!

OpenStudy (anonymous):

how 'bout this then.... \(\large \frac{sin(2x)}{3x}=\frac{2sinxcosx}{3x}=\frac{2}{3}\cdot \frac{sinx}{x}\cdot cosx \)